# Origin of effective graviton mass from spontaneous symmetry breaking of scalar field

### v1

@@ Abstract
General relativity and quantum field theory govern all fundamental physical processes, from nuclear and subatomic scales to the solar system, galactic, and cosmology. Both theories work remarkably well in their tested domains. However, there is minimal overlap between the two domains. Our work demonstrates that the non-overlapping of domains no longer holds when non-perturbative effects are considered. We introduce a novel mechanism in which the gravitons acquire effective mass due to the spontaneous symmetry breaking (SSB) of global or local symmetry of complex scalar field minimally coupled to gravity. In the flat space-time limit, we demonstrate that the resulting action for the graviton is identical to the extended Fierz-Pauli (FP) action, corresponding to the mass deformation parameter $\alpha=1/2$. The extended FP action has 6 Degrees of freedom and no ghost. For local $U(1)$ SSB, we establish that the effective graviton mass is related to the mass of the gauge boson and the electric charge of the complex scalar field. Interestingly, our results indicate that the millicharged dark matter scalar fields, generating dark photons, can produce graviton mass of the same order as the Hubble constant $(H_{0})$. Hence, we argue that the dark sector offers a natural explanation for the acceleration of the current Universe.
@@ I Introduction
The effect of graviton mass on the cosmological scales is analogous to the propagation of electromagnetic (EM) waves in plasma. In unmagnetized neutral plasma, the dispersion relation for EM waves is $\omega^{2}=k^{2}+\omega_{p}^{2}$, where $\omega_{p}$ is plasma frequency. Thus, waves with frequencies lower than $\omega_{p}$ cannot propagate in plasma. Similarly, in the case of massive gravity, high frequencies propagate in the background space-time without any attenuation. However, the low-frequency waves are attenuated, leading to a phenomenon known as degravitation [64, 65, 66]. As the CC is inherently a constant, it operates like a source with an exceedingly long wavelength [67]. Hence, it acts as a screen that serves as a high-pass filter, leading to a considerably diminished observed effective value [66].
@@
However, massive gravity theories do not naturally explain the origin of the graviton mass, leading to considerable skepticism about this approach [68, 69, 70]. Our work addresses this gap by providing a natural explanation for the origin of effective graviton mass through spontaneous symmetry breaking (SSB) of a $U(1)$ complex scalar field.
@@
SSB is a unifying concept that explains, among other things, ferromagnetism, superconductivity, Bose-Einstein condensation, and the Higgs mechanism [71, 72, 73, 74, 75, 76]. SSB arises from the interplay between nonlinearity and symmetric potentials. It is well known that the symmetry of the ground state in models with nonlinear interactions mirrors the symmetry of the external/self-interacting potential [72]. This is always true in quantum mechanics (and any linear theory) as long as the nonlinearity remains sufficiently weak. However, SSB occurs when the nonlinearity reaches a critical strength. This phenomenon is easily observed in double-well potentials [72]. SSB plays a pivotal role in the standard model of particle physics, particularly about the Higgs [77, 78, 79, 80].
@@
We demonstrate that SSB of a scalar field can lead to non-trivial infrared (IR) effects on gravity. By using the complex scalar field minimally coupled to gravity, we present a unique approach for the origin of graviton mass. The effective mass of the graviton is related to the mass of the scalar field and the self-interaction coupling constant. In the case of local U(1) SSB, we show that the graviton mass is related to the mass of the gauge boson and the electric charge of the complex scalar field. Our results show that the millicharged dark matter scalar fields that generate dark photons can lead to graviton mass of the same order as the Hubble constant $(H_{0})$. Consequently, the mechanism provides a natural explanation for the current acceleration of the universe. In the remainder of this work, we explicitly demonstrate how the graviton mass arises from the non-zero VEV of the complex scalar field in three different scenarios: Global $U(1)$ invariance, local $U(1)$ invariance and Coleman-Weinberg mechanism [81, 82].
II Effective graviton mass through global $U(1)$ SSB
@@
To keep things transparent, we start with a complex scalar field ($\Phi$) minimally coupled to gravity in a general 4-D spacetime. The action, in $\hbar=c=1$ units, is:
where $\kappa^{2}=8\pi G$. The action of a global $U(1)$ invariant complex scalar field is:
@@
where $U(|\Phi|)$ is self-interacting potential. Varying the above action (1) w.r.t the metric leads to the Einstein’s equation where the energy-momentum tensor $(T_{\mu\nu})$ is:
@@
It is easy to verify that the conservation of the stress-tensor $\nabla^{\mu}T_{\mu\nu}=0$ leads to the equations of motion. We now define a fourth-rank mixed tensor which is the metric variation of $T_{\mu\nu}$:
$\overline{\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}}(x-y)=\frac{\delta T_{\mu\nu}(x)}{\delta g_{\rho\sigma}(y)}-\frac{1}{2}g^{\rho\sigma}T_{\mu\nu}\delta^{4}(x-y)$
@@
$-\frac{1}{2}g^{\rho\sigma}T_{\mu\nu}\delta^{4}(x-y).$ (4)
Note that we can also define a related fourth-rank covariant tensor $\overline{\mathscr{O}_{\mu\nu\rho\sigma}}(x-y)$ which is obtained by the variation of the stress-tensor w.r.t the inverse metric. In this work, we will use the mixed tensor defined above. We now consider the following metric decomposition:
@@
where $\bar{g}_{\mu\nu}$ is the metric describing the background geometry w.r.t which we linearize the action (1), $h_{\mu\nu}$ is the metric perturbation and $\kappa$ carries the order of the metric perturbation. Expanding the action (1) at the quadratic level of metric fluctuation $h_{\mu\nu}$, we obtain:
$S_{h}=\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}$ (6)
-- ¯$\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}-\frac{\bar{R}}{2}\left[h_{\alpha\beta}h^{\alpha\beta}-\frac{h^{2}}{2}\right]+2\bar{R}_{\nu\sigma}\left[h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right]\Bigg{]}$
@@
$\overline{\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}}=-\frac{1}{2}\Bigg{[}\left(\delta_{\mu}^{\rho}\delta_{\nu}^{\sigma}+\delta_{\mu}^{\sigma}\delta_{\nu}^{\rho}\right)\Big{[}\bar{g}^{\alpha\beta}\partial_{\alpha}\Phi\partial_{\beta}\Phi^{*}+U(|\Phi|)\Big{]}$
@@
$~{}~{}~{}+\partial_{\nu}\Phi\partial_{\mu}\Phi^{*}-\bar{g}_{\mu\nu}\left(\bar{g}^{\alpha\beta}\partial_{\alpha}\Phi\partial_{\beta}\Phi^{*}+U(|\Phi|)\right)\Big{]}.$ (7)
[For details, see Appendices A and B.] The above action (6) is a key expression regarding which we would like to discuss the following: The last term contains the information about the variation of the stress-tensor as a function of metric. In cosmology all quantities vary with time; this is captured via Ricci scalar/tensor. At thermal equilibrium, these terms dominate over the last term. However, the contribution of the last term is significant when the Universe is out-of-equilibrium, like at SSB or phase transition [83, 84].
@@
Substituting $U(\phi_{0})$ in action (6) and using the Einstein’s equations for the background geometry, $\bar{R}=-\kappa^{2}\bar{T}$ and $\bar{R}_{\nu\sigma}=\kappa^{2}\left(\bar{T}_{\nu\sigma}-\frac{1}{2}\bar{g}_{\nu\sigma}\bar{T}\right)$ where $\bar{T}=-4U(\phi_{0})$ and $\bar{T}_{\nu\sigma}=-\bar{g}_{\nu\sigma}U(\phi_{0})$, leads to (details in Appendices A and B):
@@
$+$ ¯$\!\!\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}+\kappa^{2}U(|\phi_{0}|)\left[h_{\alpha\beta}h^{\alpha\beta}-\frac{h^{2}}{2}\right]\Big{]}.$
This is the first key result of this work, and we want to discuss the following points. First, the above action corresponds to a massive graviton. The flat spacetime limit of the above action reduces to (details in Appendix A):
@@
where,
$m={\kappa m_{\Phi}^{2}}/({2\sqrt{\lambda}})\,,\,\alpha=1/2\,.$ (13)
The above action is identical to the extended Fierz-Pauli (FP) action corresponding to massive deformation parameter $\alpha=1/2$ [86, 87]. (See section I of [88] for more details.) The above action differs from the FP action, which corresponds to $\alpha=1$ [64]. For $\alpha=1$, $h$ is traceless and $h_{\mu\nu}$ has five degrees of freedom (DoF). However, for $\alpha={1}/{2}$, $h$ satisfies the wave equation:
$(\Box-m^{2})h=0\,,$
Thus, for $\alpha=1/2$, $h_{\mu\nu}$ has six DoF. Although naively, this seems to be a ghost-like DoF [89], $\alpha={1}/{2}$ has no ghost DoF [86, 87]. In Appendices C and D, we have explicitly shown this. Thus, we conclude that the global U(1) SSB of a minimally coupled, self-interacting complex scalar field leads to an effective graviton mass which is proportional to the mass of the scalar field ($m_{\Phi}$) and inversely proportional to self-interacting coupling ($\lambda$). Second, since SSB occurs at a specific VEV, the kinetic term of the scalar field vanishes, and only the value of the potential at SSB contributes to the effective graviton mass.
Third, it is essential to note that although the global $U(1)$ symmetry breaking leads to an effective mass of gravitons, it also produces a Goldstone Boson, which would produce a long-range attractive potential. Since this is inevitable and its existence is questionable due to various observational constraints [90], this SSB phenomenon might be debatable. To overcome this, we must look for a scenario where massless modes do not occur. This naturally leads us to generalize our analysis to local $U(1)$ invariant theories.
III Effective graviton mass through local U(1) SSB

### v2

# Generation of effective massive Spin-2 fields through spontaneous symmetry breaking of scalar field

@@ Abstract
General relativity and quantum field theory are the cornerstones of our understanding of physical processes, from subatomic to cosmic scales. While both theories work remarkably well in their tested domains, they show minimal overlap. However, our research challenges this separation by revealing that non-perturbative effects bridge these distinct domains. We introduce a novel mechanism wherein, at linear order, spin-2 fields around an arbitrary background acquire effective mass due to the spontaneous symmetry breaking (SSB) of either global or local symmetry of complex scalar field minimally coupled to gravity. The action of the spin-2 field is identical to the extended Fierz-Pauli (FP) action, corresponding to the mass deformation parameter $\alpha=1/2$. We show that this occurs due to the effect of SSB on the variation of the energy-momentum tensor of the matter field, which has a dominant effect during SSB. The extended FP action has a salient feature, compared to the standard FP action: the action has 6 degrees of freedom with no ghosts. For local $U(1)$ SSB, we establish that the effective mass of spin-2 fields is related to the mass of the gauge boson and the electric charge of the complex scalar field. Interestingly, our results indicate that the millicharged dark matter scalar fields, generating dark photons, can produce a mass of spin-2 fields of the same order as the Hubble constant $(H_{0})$. Hence, we argue that the dark sector offers a natural explanation for the acceleration of the current Universe.
@@ I Introduction
The effect of mass of spin-2 fields on the cosmological scales is analogous to the propagation of electromagnetic (EM) waves in plasma. In unmagnetized neutral plasma, the dispersion relation for EM waves is $\omega^{2}=k^{2}+\omega_{p}^{2}$, where $\omega_{p}$ is plasma frequency. Thus, waves with frequencies lower than $\omega_{p}$ cannot propagate in plasma. Similarly, in the case of massive gravity, high frequencies propagate in the background space-time without any attenuation. However, the low-frequency waves are attenuated, leading to a phenomenon known as degravitation [64, 65, 66] (see Appendix D). As the CC is inherently a constant, it operates like a source with an exceedingly long wavelength [67]. Hence, it acts as a screen that serves as a high-pass filter, leading to a considerably diminished observed effective value [66].
@@
However, massive gravity theories do not naturally explain the origin of the mass of spin-2 fields, leading to considerable skepticism about this approach [68, 69, 70]. Also, the standard Fierz-Pauli action [71] corresponding to the massive spin-2 field in arbitrary curved space-time has a ghost [72] (see the discussion in Appendix C). Recently, it was shown that there exists a nonlinear completion of the Fierz-Pauli action that is free of ghost up to the fourth order [73, 74, 75]). Our work addresses this gap by providing a natural explanation for the generation of effective mass of spin-2 fields through spontaneous symmetry breaking (SSB) of a $U(1)$ complex scalar field. We also show that the extended Fierz-Pauli action does not have ghost DoF and, hence, is stable in the linear regime.
@@
SSB is a unifying concept that explains, among other things, ferromagnetism, superconductivity, Bose-Einstein condensation, and the Higgs mechanism [76, 77, 78, 79, 80, 81]. SSB arises from the interplay between nonlinearity and symmetric potentials. It is well known that the symmetry of the ground state in models with nonlinear interactions mirrors the symmetry of the external/self-interacting potential [77]. This is always true in quantum mechanics (and any linear theory) as long as the nonlinearity remains sufficiently weak. However, SSB occurs when the nonlinearity reaches a critical strength. This phenomenon is easily observed in double-well potentials [77]. SSB plays a pivotal role in the standard model of particle physics, particularly about the Higgs [82, 83, 84, 85, 86].
@@
We demonstrate that SSB of a scalar field can lead to non-trivial infrared (IR) effects on gravity. By using the complex scalar field minimally coupled to gravity, we present a unique approach for the origin of mass of spin-2 fields. The effective mass of the spin-2 fields is related to the mass of the scalar field and the self-interaction coupling constant. In the case of local U(1) SSB, we show that the mass of spin-2 fields is related to the mass of the gauge boson and the electric charge of the complex scalar field. Our results show that the millicharged dark matter scalar fields that generate dark photons can lead to graviton mass of the same order as the Hubble constant $(H_{0})$. Consequently, the mechanism provides a natural explanation for the current acceleration of the universe. In the remainder of this work, we explicitly demonstrate how the mass of spin-2 fields arises from the non-zero VEV of the complex scalar field in three different scenarios: Global $U(1)$ invariance, local $U(1)$ invariance and Coleman-Weinberg mechanism [87, 88].
The rest of this work is organized as follows: In Sec. (II), we discuss the setup and define the fourth-order tensor ${\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}}$. As a first application of our work, in Sec. (III), we obtain the effective mass for the spin-2 field via global U(1) SSB. In Sec. (IV), we delve deeper into our analysis and demonstrate that effective massive spin-2 is generated via local U(1) SSB. Sec. (V) discusses the key results and possible implications in the early Universe cosmology. Appendices (A) - (D) provide details of the calculations in the main text. In this work, we set $\hbar=c=1$ and $\kappa^{2}=8\pi G$.
II The Setup
@@
To keep things transparent, we start with a complex scalar field ($\Phi$) minimally coupled to gravity in a general 4-D spacetime. The action is:
$S=\frac{1}{2\kappa^{2}}\int\!\!d^{4}x\sqrt{-g}R+S_{\Phi}(\Phi,\partial_{\mu}\Phi)\,.$ (1)
The action of a global $U(1)$ invariant complex scalar field is:
@@
where $U(|\Phi|)$ is self-interacting potential. Varying the above action (1) w.r.t the metric leads to Einstein’s equations ($G_{\mu\nu}=\kappa^{2}\,T_{\mu\nu}$), where the energy-momentum tensor $(T_{\mu\nu})$ is:
@@
For a generic complex scalar field $\Phi(t,\bf{x})$, there are ten components of the energy-momentum tensor $T_{\mu\nu}$. On the other hand, the conservation of the energy-momentum tensor leads to 6 independent components of the energy-momentum tensor. In a similar manner, $\nabla^{\mu}G_{\mu\nu}=0$ leads to 6 independent components of Einstein’s tensor in general. As a result, the number of independent components for a generic complex scalar field remains the same on both sides of Einstein’s field equations. This is consistent with the number of independent metric components in arbitrary 4-D spacetime being 6. It is easy to verify that the conservation of the stress-tensor $\nabla^{\mu}T_{\mu\nu}=0$ leads to the equations of motion. We now define a fourth-rank mixed tensor which is the metric variation of $T_{\mu\nu}$:
${\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}}$ $=$ $\frac{\delta T_{\mu\nu}(x)}{\delta g_{\rho\sigma}(y)}-\frac{1}{2}g^{\rho\sigma}T_{\mu\nu}\delta^{4}(x-y)$
@@
-- $\frac{1}{2}g^{\rho\sigma}T_{\mu\nu}\delta^{4}(x-y)$
Note that we can also define a related fourth-rank covariant tensor ${\mathscr{O}_{\mu\nu\rho\sigma}}(x-y)$ which is obtained by the variation of the stress-tensor w.r.t the inverse metric. In this work, we will use the mixed tensor defined above. It is important to note that, like $T_{\mu\nu}$, the fourth rank ${\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}}$ is defined for any arbitrary space-time. Note that the presence of the Dirac delta function ensures that the operator is local, ultimately resulting in a local action for the spin-2 field.
Einstein’s field equations are highly non-linear, which restricts the possibility of finding solutions to only a few highly symmetric cases. The perturbative approach is essential to understand the nature of the exact solutions beyond the symmetric cases. However, to conduct a perturbative expansion, it is necessary to establish a background — a manifold representing the average (Universe) or background space-time. The process of averaging in GR is ambiguous due to the involvement of complex non-linear interactions. The influence of these non-linearities on the overall expansion is referred to as backreaction, which remains significant even at late times [89, 90].
In our approach, we will refrain from making any assumptions about the background metric ($\bar{g}_{\mu\nu}$) for two reasons: First, we aim to capture the influence of the non-linear potential on perturbations, particularly those arising from spontaneous symmetry breaking (SSB). In other words, we want to examine the contribution of the SSB potential to the background itself. Therefore, our analysis requires a non-flat background metric, denoted by $\bar{g}_{\mu\nu}$. Second, the potential $U(|\Phi|)$ that we will consider is highly non-linear, and because the scalar field is dynamic, it is not possible to impose any specific symmetry on the background metric.
As we show in the rest of this work, despite the unknown exact form of the background metric, our approach adeptly captures the dynamics of SSB on the spin-2 fields. Based on this, we now consider the following metric decomposition:
@@
where $\bar{g}_{\mu\nu}$ is the metric describing the background geometry w.r.t which we linearize the action (1), $h_{\mu\nu}$ is the metric perturbation and $\kappa$ carries the order of the metric perturbation [^1] By linearizing, we refer to the decomposition presented in the above equation where $2\kappa|h_{\mu\nu}|\ll|\bar{g}_{\mu\nu}|$. This essentially implies that $h_{\mu\nu}$ can be treated as a metric perturbation with respect to the background metric $\bar{g}_{\mu\nu}$, which satisfies the background Einstein field equations.. Since the background metric is arbitrary, it has 6 independent components. Hence, the metric perturbations $h_{\mu\nu}$ also have six independent components.
Expanding the action (1) at the quadratic level of metric fluctuation $h_{\mu\nu}$ (details in Appendix A), we obtain:
$+$ ¯$\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}-\frac{\bar{R}}{2}\left[h_{\alpha\beta}h^{\alpha\beta}-\frac{h^{2}}{2}\right]$
$+$ $2\bar{R}_{\nu\sigma}\left[h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right]\Bigg{]}$
@@
$\overline{\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}}=-\frac{1}{2}\Bigg{[}\left(\delta_{\mu}^{\rho}\delta_{\nu}^{\sigma}+\delta_{\mu}^{\sigma}\delta_{\nu}^{\rho}\right)\Big{[}\bar{g}^{\alpha\beta}\partial_{\alpha}\Phi\partial_{\beta}\Phi^{*}+U(|\bar{\Phi}|)\Big{]}$
@@
$~{}~{}~{}+\partial_{\nu}\Phi\partial_{\mu}\Phi^{*}-\bar{g}_{\mu\nu}\left(\bar{g}^{\alpha\beta}\partial_{\alpha}\Phi\partial_{\beta}\Phi^{*}+U(|\bar{\Phi}|)\right)\Big{]}\,,$ (7)
where $\overline{\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}}$ is the fourth-rank tensor (II) obtained w.r.t. the arbitrary background metric $\bar{g}_{\mu\nu}$. The above action (6) is a key expression regarding which we would like to discuss the following: The last term contains the information about the variation of the stress-tensor as a function of metric. In cosmology all quantities vary with time; this is captured via Ricci scalar/tensor. At thermal equilibrium, these terms dominate over the last term. This will be explicitly shown in Sec. (IV), where we consider a local U(1) SSB of a complex scalar field. However, the contribution of the last term is significant when the Universe is out-of-equilibrium, like at SSB or phase transition [91, 92].
As mentioned earlier, GR is highly non-linear, and obtaining exact solutions is possible only for specific symmetric space-times. Given this, perturbation theory in GR can be approached in two ways: The first approach does not assume any symmetry of the background metric and only requires that it be a solution to Einstein’s field equations. Alternatively, one can assume a specific form for the background metric and utilize its symmetries to derive the perturbation equations. The second approach is more widely used in black hole physics and cosmology [93, 94, 95].
In this work, we use the first approach and the fourth-rank tensor (II) contributes to the backreaction of the stress-energy tensor onto the background space-time. The second approach involves extracting the constant $U_{0}$ from the symmetry-breaking field potential $U(|\Phi|)$ at the tree-level action. The contribution to the mass of spin-2 field coming from the field $\Phi$ after SSB will vanish since we have already extracted the non-zero minima value $U_{0}$ from the field potential.
III Effective massive spin-2 fields through global $U(1)$ SSB
@@
Substituting $U(\phi_{0})$ in action (6) and using the Einstein’s equations for the background geometry, $\bar{R}=-\kappa^{2}\bar{T}$ and $\bar{R}_{\nu\sigma}=\kappa^{2}\left(\bar{T}_{\nu\sigma}-\frac{1}{2}\bar{g}_{\nu\sigma}\bar{T}\right)$ where $\bar{T}=-4U(\phi_{0})$ and $\bar{T}_{\nu\sigma}=-\bar{g}_{\nu\sigma}U(\phi_{0})$, leads to (details in Appendix A):
@@
$+$ ¯$\!\!\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}-\frac{1}{2}m^{2}(h_{\alpha\beta}h^{\alpha\beta}-\alpha h^{2})\Big{]}\,$
where,
$m={\kappa m_{\Phi}^{2}}/({2\sqrt{\lambda}})\,,\,\alpha=1/2\,.$ (12)
This is the first key result of this work, and we want to discuss the following points. First, the above action corresponds to a massive spin-2 field in arbitrary curved spacetime and is known as the generalized Fierz-Pauli action with $\alpha$ being massive deformation parameter (see Appendix B for more details). It is important to note that the Fierz-Pauli action is a special case of the above action where $\alpha=1$.
Second, while the Fierz-Pauli action in Minkowski spacetime has only five dynamical polarizations, it has six in arbitrary curved spacetime. The additional polarization is a ghost [72] (see the discussion in Appendix C). As can be seen from Eq. (6), the Fierz-Pauli theory does not have any ghost DOF in Einstein spaces, defined by the relation $R_{\mu\nu}=\Lambda g_{\mu\nu}$. However, this can not extended to arbitrary background as the action depends on the background Ricci tensor and Ricci scalar [96, 97, 98, 99, 100]. Recently, nonlinear extension of massive gravity gains much attention in order to take into account the IR effect within GR (see, for instance, [73, 74, 75]). Here, as shown in Appendices C and D, the extended Fierz-Pauli action does not have ghost DoF and hence is stable in the linear regime.
The flat spacetime limit of the above action reduces to (details in Appendix B.2):
@@
The above action is identical to the extended Fierz-Pauli (FP) action corresponding to massive deformation parameter $\alpha=1/2$ [101, 102]. (See Appendix B for more details.) The above action differs from the FP action, corresponding to $\alpha=1$ [64]. For $\alpha=1$, $h$ is traceless, and $h_{\mu\nu}$ has five degrees of freedom (DoF). However, for $\alpha={1}/{2}$, $h$ satisfies the wave equation:
$(\Box-m^{2})h=0.$ (14)
Thus, for $\alpha=1/2$, $h_{\mu\nu}$ has six DoF. Although naively, this seems to be a ghost-like DoF [103], $\alpha={1}/{2}$ has no ghost DoF [101, 102]. In Appendices B and C, we have explicitly shown that the extended Fierz-Pauli action has no ghosts. Hence, the theory is well-defined in arbitrary curved spacetime. It is important here to note that since the background metric $\bar{g}_{\mu\nu}$ is chosen to be arbitrary, it has six independent components, which also implies that the resultant metric perturbation $h_{\mu\nu}$ has 6 DoF which is mathematically consistent with our analysis. On the other hand, it is also important to note that in asymptotic spatial infinity, the DoF of metric perturbation reduces to 2 as the energy-momentum vanishes there identically. As a result, there would not be any SSB in the matter sector. Thus, we conclude that the global U(1) SSB of a minimally coupled, self-interacting complex scalar field leads to an effective mass of spin-2 fields which is proportional to the mass of the scalar field ($m_{\Phi}$) and inversely proportional to self-interacting coupling ($\lambda$).
Third, since SSB occurs at a specific VEV, the kinetic term of the scalar field vanishes, and only the value of the potential at SSB contributes to the effective mass of spin-2 fields. Fourth, it is essential to note that although the global $U(1)$ symmetry breaking leads to an effective mass of spin-2 fields, it also produces a Goldstone Boson, which would produce a long-range attractive potential. Since this is inevitable and its existence is questionable due to various observational constraints [104], this SSB phenomenon might be debatable. To overcome this, we must look for a scenario where massless modes do not occur. This naturally leads us to generalize our analysis to local $U(1)$ invariant theories.
Lastly, as previously noted, the non-linear nature of Einstein’s field equations limits the availability of exact solutions, making a perturbative approach essential. Although the averaging process in GR is ambiguous due to complex non-linear interactions, our method effectively captures the influence of the non-linear potential on perturbations, particularly as it relates to SSB. The traditional perturbative approach used in black hole physics and cosmology [93, 94, 95] typically assumes a background geometry ab initio. As mentioned earlier, our approach is different from the standard perturbation theory. The standard perturbation theory involves isolating the constant $U_{0}$ from the field potential at the tree-level action, followed by decomposing the metric into a background component and perturbations. This ultimately yields the same expression for the mass of the tensor perturbations as obtained earlier in this section. However, this approach carries a limitation: the background geometry is assumed to be either de Sitter or anti-de Sitter, solely based on $U_{0}$. In contrast, the approach used in this work does not make any specific assumptions about the background metric, other than that it satisfies the background Einstein field equations. It is particularly effective in capturing the influence of the non-linear potential due to SSB, even when the exact form of the background metric remains undetermined.
IV Effective massive spin-2 fields through local U(1) SSB